OpenStudy (anonymous):

Solve 6 and two-ninths minus 4 and one-third. Show your work.

hero (hero):

:(

hero (hero):

I would have shown you exactly how to do this

hero (hero):

The method is slightly different, but not that difficult

hero (hero):

\(6\frac{2}{9} - 4\frac{1}{3}\) (\(6-4)(\frac{2}{9} - \frac{1}{3})\) \((6-4)(\frac{2}{9} - \frac{3}{9})\) \((5-4)(\frac{11}{9} - \frac{3}{9})\) \(1\frac{8}{9}\)

hero (hero):

If you want, I can actually show you how to get that.